For the function y =[(x-a)/(a^2 -b^2)]
and the initial value of the independent variable x0 = a - b. At what terminal value of x the increment ?y = 1/(a-b) .

Grade:11

1 Answers

bharat bajaj IIT Delhi
10 years ago
y = 1/a-b = x-a / (a+b)(a-b)
Hence, 1 = x-1/a+b
a+ b = x - 1
Hence, terminal value of x = a+b + 1
Thanks
Bharat Bajaj
IIT Delhi
askiitians faculty